OpenStudy (anonymous):

verify the identify cos(2A)=1-2sin^2(A) using Cos (A+B)=cosA CosB- SinA SinB

OpenStudy (gorv):

\[\cos2A=\cos(A+A)\]

OpenStudy (gorv):

can u do it further??

OpenStudy (gorv):

expand it

OpenStudy (gorv):

cosA*cosA-sinA*sinA

OpenStudy (gorv):

cos^2A-sin^2A

OpenStudy (gorv):

cos^2A=1-sin^2A

OpenStudy (gorv):

plug it in place of cos^2A

OpenStudy (anonymous):

you lost me at cosA*cosA-sinA*sinA

OpenStudy (anonymous):

sum of angles identity is cos (A+B) =cosAcosB-sinAsinB in this case B=A so just sub it in gurl

OpenStudy (anonymous):

cos(A+A)=cos ^2 A-sin ^2 A=1-2sin ^2 A since, sin ^2 A+ cos ^2 A=1

OpenStudy (anonymous):

@sierralake